# Time Resolved Absorption of Six Chemical Species With MAROON-X Points to a Strong Drag in the Ultra Hot Jupiter TOI-1518 b

### v2

@@ Abstract
Context. Wind dynamics play a pivotal role in governing transport processes within planetary atmospheres, influencing atmospheric chemistry, cloud formation, and the overall energy budget. Understanding the strength and patterns of winds is crucial for comprehensive insights into the physics of ultra-hot Jupiter atmospheres. Current research has proposed two contrasting mechanisms that limit wind speeds in these atmospheres, each predicting a different scaling of wind speed with planet temperature. However, the sparse nature of existing observations hinders the determination of population trends and the validation of these proposed mechanisms.
Aims. This study focuses on unraveling the wind dynamics and the chemical composition in the atmosphere of the ultra-hot Jupiter TOI-1518 b
Methods. Two transit observations using the high-resolution ( 000), optical (spectral coverage between 490 and 920 nm) spectrograph MAROON-X were obtained and analyzed to explore the chemical composition and wind dynamics using the cross-correlation techniques, global circulating models, and atmospheric retrieval.
Results. We report the detection of 14 species in the atmosphere of TOI-1518 b through cross-correlation analysis. Additionally, we measure the time-varying cross-correlation trails for 6 different species, compare them with predictions from General Circulation Models (GCM) and conclude that a strong drag is present in TOI-1518b’s atmosphere (). The ionized species require stronger drags than neutral species, likely due to the increased magnetic effects in the upper atmosphere. Furthermore, we detect vanadium oxide (VO) using the most up-to-date line list. This result is promising in detecting VO in other systems where inaccuracies in previous line lists have hindered detection. We use a retrieval analysis to further characterize the abundances of the different species detected. Chromium, magnesium, vanadium, and titanium are found with lower than solar abundances, possibly due to the ionization of these elements or to their incorporation into TiO and VO.
@@ 1 Introduction
When using low to moderate spectral resolution (), like that provided by a space telescope such as JWST, the spectra captured contain a mixture of information from various parts of the atmosphere. Each part of the atmosphere has different properties, such as temperature or chemical composition (Espinoza et al., 2024). This information mixture could lead to misleading or biased inferences about the different properties derived from the data (Feng et al., 2015; Line & Parmentier, 2016).
@@
For the first time in 2010, high-resolution () spectroscopy was used to characterize the atmosphere of a transiting ultra-hot Jupiter by resolving individual molecular lines using CRIRES at VLT (Snellen et al., 2010). During the transit, the Doppler shifts caused by the planet’s rotation and the atmospheric winds allow lines formed in different parts of the planetary atmosphere to be spectroscopically separated (see Nortmann et al. 2024 for an extreme example). Recent ESPRESSO observations at VLT showed that the iron absorption lines of WASP-76b and WASP-121b, two canonical Ultra-Hot Jupiters (UHJs), are progressively blueshifted during the transit (Ehrenreich et al., 2020; Borsa et al., 2021).
@@
While two competing scenarios have been suggested to explain this behavior, the precise physical mechanism remains elusive.
@@
The signal could result from a hot, puffy evening terminator and a cool, compact morning terminator, whereby the blueshifting winds and rotation of the hot evening terminator dominate the absorption signal due to its larger-scale height (Wardenier et al., 2021). Alternatively, it has been shown that 3D models with opaque clouds can also reproduce the observed signal (Savel et al., 2022), whereby the cloud deck “blocks” the absorption features on the morning terminator. Other studies highlights the richness of WASP-76b with the detection of multiple species with different shift (Kesseli et al., 2022).
To get more insights into this wind speed problem, we observed the UHJ TOI-1518 b. This planet is orbiting a fast-rotating F0-type star of 7300K. Table 1 summarizes the physical properties of TOI-1518 and TOI-1518 b. This planet (T = 2498 K) is ideal for a direct comparison with the well-studied UHJs WASP-76 b (T = 2228 K) and WASP-121 b (T = 2720 K) with iron previously detected by Cabot et al. (2021).
@@
In this paper, we present the observations of two transits of TOI-1518 b observed with MAROON-X at the Gemini-North Observatory. After presenting the observations and the data reduction in Sect. 2, we present the chemical information we obtained thanks to Cross-Correlation techniques in Sect. 3. Then, we compare the iron trail detected with global circulating models (GCMs) to explore the wind dynamics of the planet in Sect. 4. We finally present a retrieval analysis in Sect. 5, hinting at the different abundances of the species detected in the atmosphere of TOI-1518 b.
| Stellar parameters | Value | Ref |
| Stellar radius | [1] | |
| Effective temperature | [1] | |
| Metallicity [Fe/H] | [1] | |
| Rotational velocity | [1] | |
| Spectral type | F0 | [1] |
| Planetary parameters | Value | Ref |
| Planet mass | [1] | |
| Planet radius | [1] | |
| Equilibrium temperature | [1] | |
| System parameters | Value | Ref |
| Orbital period | [1] | |
| Mid-transit Time | [1] | |
| Orbital Inclination | [1] | |
| Semi-major axis | [1] | |
| Systemic Velocity () | [1] | |
| Projected orbital velocity () | ||
| Impact Parameter (b) | [1] |
@@ 2 Observations and Data reduction
Two transits of the UHJ TOI-1518 b were observed with MAROON-X, a high-resolution () optical (spectral coverage between 490 and 920 nm) spectrograph at the 8.1-m Gemini-North observatory in Hawaii. Recent observations showed the capacity of MAROON-X to characterize UHJs by detecting ions and volatile and refractory elements. It has also allowed the description of the trail of the atmospheric signal (Pelletier et al., 2023; Prinoth et al., 2023) and even to study some strong lines such as Ca+ triplet (Prinoth et al., 2024). The observations were taken on 2022-08-13 and 2023-10-19 (program ID GN-2022B-Q-128 and GN-2023B-Q-127, PI: Parmentier). A summary of the observations is given in Table 2.
@@ 2.1 Data Quality
MAROON-X is divided into two detectors, one ”blue” covering wavelengths ranging from 490 to 678 nm. The second one, ”red,” covers wavelengths ranging from 640 to 920 nm. To ensure complete coverage of transit events, the observation protocol includes baseline measurements both pre- and post-transit. The exposure time for the red detector was slightly lower than for the blue arm (220s vs 260s) during the first observation. For the second observation, the exposure time for both detectors was set up to 220 sec. Each order of the red detector comprises 4036 pixels, while the blue one has only 3954 pixels. There are 28 spectral orders for the red detector and 33 for the blue one. Fig 1 presents the signal-to-noise ratio (S/N) and the airmass as a function of the observation frames for both transits. For both nights, the S/N was at least above 110, with a slightly better S/N for the second night (155 min and 210 max). During the second night, the blue arm seemed less performant than the red arm, but this is because the blue arm’s exposure time was higher during the first transit. Conditions of the second transit (average humidity = 7%) were so good that both arms still had better S/N than during the first transit (average humidity =24%).
@@
One main limitation with high-resolution transmission spectroscopy from ground-based observations is our atmosphere and stellar signals. Planetary signals are much fainter but change over time because of the rapid Doppler acceleration, inducing shifts of many tens of km.s-1 to the planet spectrum over the transit duration. The telluric lines stay constant, and the positions of the stellar lines vary in order of hundreds of m.s-1. This is why we can disentangle the planetary signal from the others. We applied different reduction steps on the data following the sequence described by Pelletier et al. (2023) and summarized below:
@@
The in-transit data are divided by a master stellar spectrum made with the out-of-transit data.
A principal component analysis (PCA) approach removes the telluric signal and residuals from the stellar correction. We selected three number of principal components to remove from each species. Higher number of components would start affecting the planetary atmospheric signal. The Ca+ triplet is so strong that no PCA is needed to detect it, as shown in Fig. 2. Both panels show two orders of MAROON-X red detectors where the lines of the Ca+ triplet are located. The top panel shows the data after pipeline extraction with the strong Ca+ stellar lines. The lower panel shows the residuals of the data reduction that preceded the PCA. The strong stellar Ca+ lines are removed using the out-of-transit data, leaving the planetary signal with the Doppler shadow effect (see sub-Section 3.3). After this step, the cleaned data are ready for Cross-Correlation to study the atmospheric composition of TOI-1518 b.
@@ 3.1 Template spectra for cross-correlation
The Cross-Correlation method is needed to detect the faint planetary lines in the residuals obtained after PCA. We cannot detect most of them directly except for a few individual lines (such as Ca+). Fortunately, atoms and molecules have so many spectral lines that combining the signal with CCF boosts the S/N, allowing a precise detection of the planetary signal. A template is needed to combine these lines. Synthetic spectra of TOI-1518 b have been generated with PetitRadtrans (Mollière et al., 2019) and FastChem (Stock et al., 2022), which respect equilibrium chemistry.
The spectra for individual molecules were produced at a resolution of over the 400 to 1000 nm wavelength range. The input parameters are those of the TOI-1518 system as shown in Table. 1. Collision induced absorption (CIA) cross-sections for the H2-H2 and H2-He pairs are used to generate a continuum opacity.
Subsequently, these spectra have been interpolated onto the MAROON-X wavelength grid and convolved to match the instrumental resolution.
The spectra were also convolved with the planetary rotation rate. These spectra are shown in Fig. 3. The species selected in this study are based on those previously detected in recent MAROON-X publications (Prinoth et al., 2023; Pelletier et al., 2023).
Line list from Kurucz database were used for all atoms and ions (Kurucz, 2017). For TiO, we used TOTO line list (McKemmish et al., 2019). For VO, both the HyVO line list (Bowesman et al., 2024) and the VOmyt line list (McKemmish et al., 2016) outputs were compared in Sect.3.2 and 3.4.
Alkaline metals and ions show individual solid lines, while other metals show line forests. Most of the signals except for the Ca+ triplet are stronger in the blue part of the spectrum, which corresponds to the blue detector of MAROON-X. We decided to analyze each detector individually as if it were two different transits and then to sum every - map or CCF map.
@@ 3.2 Species detected in the - maps
The cross-correlation maps are converted into velocity-velocity maps (- diagrams) by shifting them towards the expected rest-frame of the planet ( = 218 6.79 km.s-1, = 0.17 km.s-1), assuming values of projected orbital velocity between 0 and 400 km.s-1 in steps of 1 km. s-1. Fig. 4 shows fourteen clear detections obtained thanks to the CCF. The white cross in each plot is the position of the expected signal from the planet if the planetary atmosphere is considered static and the planet has a circular orbit. The signal is in the rest frame of the planet due to the previous correction of the reduction process from the systemic and barycentric velocity. The expected is calculated as follows:
with , where is the period, is the semi-major axis of TOI-1518 b and the inclination of the system. All these parameters are given in Table 1. To compute our - plots, we selected a range of orbital velocity () from 0 to 400 km.s-1 and a range of rest-frame velocity () from -150 to 150 km.s-1 with steps of 1 km.s-1 for each. We then integrate each point of the CCF maps previously obtained (see Fig. 4) following the slope determined by the orbital velocity at the rest frame position determined by . Fourteen species are detected with an S/N 4.5. The noise level is calculated in a region far from the central peak at 75 km.s-1 where no signal of the planet or Rossiter-McLaughlin (RM) residuals is expected. The parameters of the best Gaussian fits of the maps are presented in Table 3.
@@
| Ca+ | -84.8 0.5 | 21.1 0.5 | 4.9 0.1 | 36.9 0.2 |
| Fe+ | -17.8 0.1 | 12.3 1.1 | -4.0 0.5 | 11.4 1.2 |
| Ti+ | -26.7 0.1 | 7.5 1.1 | -3.4 0.6 | 9.1 1.4 |
| Ba+ | -26.6 0.1 | 11.4 -0.6 | -5.2 0.3 | 13.7 0.8 |
| Ca | -22.1 0.1 | 8.1 0.9 | -3.7 0.6 | 10.1 1.4 |
| Fe | -32.9 0.2 | 17.9 1.7 | -3.8 0.5 | 11.1 1.2 |
| Ti | -15.2 0.4 | 10.1 0.7 | -4.3 0.3 | 9.7 0.8 |
| Si | -39.9 0.2 | 5.9 0.9 | -1.4 0.8 | 9.7 1.8 |
| Mg | -0.8 0.8 | 8.1 0.8 | -3.1 0.6 | 11.7 1.4 |
| V | -58.1 0.6 | 5.8 0.5 | -3.1 0.5 | 10.2 1.1 |
| Mn | -16.2 0.6 | 4.0 0.2 | -3.5 0.2 | 12.7 0.6 |
| Cr | -19.5 0.1 | 4.0 0.4 | -1.3 0.8 | 12.5 1.9 |
| Na | +13.4 2.0 | 7.9 1.0 | -1.3 1.2 | 18.3 3.3 |
| VO | -43.5 0.2 | 4.9 0.2 | -12.1 0.2 | 10.5 0.5 |
| All | -31.2 0.1 | 21.1 2.6 | -2.9 0.7 | 11.6 1.7 |
The of other species of interested detected in other UHJs are presented in Appendix (see Fig 17). These ones show positive correlation near the expected orbital position and may warrant follow-up observations. For some species, such as iron, calcium, magnesium, or titanium, the signal is blueshifted compared to the expected velocity (Vres = km.s-1). The orbital velocity is also lower than expected (km.s-1).

### v3

@@ Abstract
Context. Wind dynamics play a pivotal role in governing transport processes within planetary atmospheres, influencing atmospheric chemistry, cloud formation, and the overall energy budget. Understanding the strength and patterns of winds is crucial for comprehensive insights into the physics of ultra-hot Jupiter atmospheres. Current research has proposed different mechanisms that limit wind speeds in these atmospheres.
Aims. This study focuses on unraveling the wind dynamics and the chemical composition in the atmosphere of the ultra-hot Jupiter TOI-1518 b.
Methods. Two transit observations using the high-resolution ( 000), optical (spectral coverage between 490 and 920 nm) spectrograph MAROON-X were obtained and analyzed to explore the chemical composition and wind dynamics using the cross-correlation techniques, global circulation models, and atmospheric retrieval.
Results. We report the detection of 14 species in the atmosphere of TOI-1518 b through cross-correlation analysis. Vanadium Oxide was detected only with the new HyVO line list whereas Titanium Oxide was not detected. Additionally, we measure the time-varying cross-correlation trails for 6 different species, compare them with predictions from Global Circulation Models (GCM) and conclude that a strong drag is slowing the winds in TOI-1518 b’s atmosphere (). We find that the trails are species dependent. favors stronger drag than Fe, which we interpret as a sign of magnetic effects being responsible for the observed strong drag. Furthermore, we show that probes layers above the Roche lobe, leading to a qualitatively different trail than the other species. Finally, We use a retrieval analysis to further characterize the abundances of the different species detected. That analysis is refined thanks to the updated planetary mass of MJup we derived from the radial-velocity detection we also report here using SOPHIE data. We mesure an abundance of iron of corresponding to 0.07 to 1.62 solar enrichment. For the other elements, the retrievals appear to be biased, probably due to the different Kp/Vsys shifts between iron and the other elements, which we demonstrate for the case of .
@@ 1 Introduction
When using low to moderate spectral resolution (), (e.g. JWST), the observed spectra contains a mixture of information from various parts of the atmosphere. Because each part of the atmosphere has different properties, such as temperature or chemical composition (Espinoza et al., 2024), this can lead to misleading or biased inferences about the different properties derived from the data (Feng et al., 2015; Line & Parmentier, 2016).
@@
For the first time in 2010, high-resolution () spectroscopy was used to characterize the atmosphere of a transiting hot Jupiter by resolving individual molecular lines using CRIRES at VLT (Snellen et al., 2010). During the transit, the Doppler shifts caused by the planet’s rotation and the atmospheric winds allow lines formed in different parts of the planetary atmosphere to be spectroscopically separated (e.g. Nortmann et al. 2024). Recent ESPRESSO observations at VLT showed that the iron absorption lines of WASP-76b and WASP-121b, two canonical Ultra-Hot Jupiters (UHJs), are progressively blueshifted during the transit (Ehrenreich et al., 2020; Borsa et al., 2021).
@@
While different scenarios have been suggested to explain this behavior, the precise physical mechanism remains elusive.
@@
The signal could result from a hot, puffy evening terminator and a cool, compact morning terminator, whereby the blueshifting winds and rotation of the hot evening terminator dominate the absorption signal due to its larger-scale height (Wardenier et al., 2021). Alternatively, it has been shown that 3D models with opaque clouds can also reproduce the observed signal (Savel et al., 2022), whereby the cloud deck “blocks” the absorption features on the morning terminator. Other studies highlights the richness of WASP-76b with the detection of multiple species with different shift (Kesseli et al., 2022). Finally, as shown by Beltz et al. (2023), magnetic effects may also contribute to the observed Doppler shift in high-temperature targets.
Here we present two transit observations of TOI-1518 b with MAROON-X at the Gemini-North Observatory. As show in Table 2, TOI-1518 b, with an equilibrium temperature of (T = 2546 K) sits in-between the well studied UHJs WASP-76 b (T = 2228 K) and WASP-121 b (T = 2720 K) with iron previously detected by Cabot et al. (2021).
@@
After presenting the observations and the data reduction in Sect. 2, we present the chemical information we obtained thanks to Cross-Correlation techniques in Sect. 3. Then, we compare the iron trail detected with global circulation models (GCMs) to explore the wind dynamics of the planet in Sect. 4. We finally present a retrieval analysis in Sect. 5, hinting at the different abundances of the species detected in the atmosphere of TOI-1518 b.
@@ 2 Observations and Data reduction
We observed two transits of the UHJ TOI-1518 b with MAROON-X, a high-resolution () optical (spectral coverage between 490 and 920 nm) spectrograph at the 8.1-m Gemini-North observatory in Hawaii. Recent observations showed the capacity of MAROON-X to characterize UHJs by detecting ions and volatile and refractory elements. It has also allowed the description of the time variation of the atmospheric signal during transit (also known as the ”trail”) (Pelletier et al., 2023; Prinoth et al., 2023) and even to study some strong lines such as triplet (Prinoth et al., 2024). The observations were taken on 2022-08-13 and 2023-10-19 (program ID GN-2022B-Q-128 and GN-2023B-Q-127, PI: Parmentier). A summary of the observations is given in Table 1.
@@ 2.1 Data Quality
MAROON-X is divided into two detectors, one ”blue” covering wavelengths ranging from 490 to 678 nm. The second one, ”red,” covers wavelengths ranging from 640 to 920 nm. To ensure complete coverage of transit events, the observations include 30min baseline measurements both pre- and post-transit. The exposure time for the red detector was slightly lower than for the blue arm (220s vs 260s) during the first observation. For the second observation, the exposure time for both detectors was set up to 220 sec. Each order of the red detector comprises 4036 pixels, while the blue one has only 3954 pixels. There are 28 spectral orders for the red detector and 33 for the blue one. Fig 1 presents the signal-to-noise ratio (S/N) and the airmass as a function of the observation frames for both transits. For both nights, the S/N was was always above 110, with a slightly better S/N for the second night (155 min and 210 max). During the second night, the blue arm looked less performant than the red arm because the blue arm’s exposure time was higher during the first transit. Conditions of the second transit (average humidity = 7% and lower airmass) were better and so both arms still have better S/N than during the first transit (average humidity =24%).
@@
One main limitation with high-resolution transmission spectroscopy from ground-based observations is the Earth’s atmosphere and stellar signals. Planetary signals are much fainter but change over time because of the rapid Doppler acceleration, inducing shifts of many tens of km.s-1 to the planet spectrum over the transit duration. The telluric lines, however, stay constant, and the positions of the stellar lines vary only on the order of hundreds of m.s-1. This allows us to distinguish the planetary signal from the other two. We applied different reduction steps on the data following the sequence described by Pelletier et al. (2023) and summarized below:
@@
The in-transit data are divided by a master stellar spectrum made with the averaged out-of-transit data.
A principal component analysis (PCA) approach removes the telluric signal and residuals from the stellar correction. For most species we removed the first three principal components and verified that this was not significantly affecting our signal (see Fig. 24). This is not the case for the lines, which dominate the observed spectrum and are thus too affected by the PCA. For these we decided to not use any PCA correction.
We show in Fig. 2 two orders of MAROON-X red detectors before (top) and after (bottom) the division by the master out stellar spectrum. The strong stellar lines are efficiently removed, leaving apparent the planetary lines and the Doppler shadow effect (see sub-Section 3.4). In this wavelength range, the planetary lines are clearly visible, even without PCA corrections.
@@ 3 Cross-Correlation analysis
3.1 TOI-1518 system
| Stellar parameters | Value |
| Stellar radius | |
| Effective temperature | |
| Metallicity [Fe/H] | |
| Rotational velocity | |
| Spectral type | F0 |
| Planetary parameters | Value |
| Planet mass | |
| Planet radius | |
| Equilibrium temperature | |
| System parameters | Value |
| Orbital period | |
| Mid-transit Time | |
| Orbital Inclination | |
| Semi-major axis | |
| Systemic Velocity () | |
| Projected orbital velocity () | |
| Impact Parameter (b) |
Notes: : derived from Eq.1.
TOI-1518 b is an ultra-hot Jupiter that was discovered by Cabot et al. (2021). It is a 1.875 R planet with an equilibrium temperature for zero albedo of 2546 K. The planet is misaligned, with an impact parameter of 0.9 (see Table.2), meaning that it doesn’t cross the zero-velocity planet of the star. The planet is precessing and the impact parameter has been shown to vary through time (Watanabe et al., 2024). Between our observations a change of 0.3 was expected, which is too small to affect our observations(from b = 0.91497 in 2019 to b= 0.8797 in 2022).The geometry of the system is presented in Fig. 18.
The planet orbit a 7300 K F-type star, which is fast rotating. The fast rotation rate from the star impeded Cabot et al. (2021) to detect the radial velocity motion due to the orbit of the planet. As such, Cabot et al. (2021) were only able to put an upper limit to the planetary mass. We also report in Appendix A the first significant detection of TOI-1518 b in radial velocity, using the SOPHIE spectrograph. This allows us to even more validate the planetary nature of the transits and to better characterize the system’s parameters, in particular the planetary mass and the semi-major axis. SOPHIE measurements allow us to pin-down the mass of TOI-1518 b to 1.83 M (see Appendix A for details). Additionally, we analyzed 56 transits of TOI-1518 b captured by TESS, compared to only 24 analyzed in Cabot et al. (2021). This increased data set has enabled us to refine the planetary ephemerides, including the semi-major axis. The improved parameters lead to a planetary semi-amplitude velocity Kp=207.32, which is 10 km/s lower (1.5 sigma lower) than in Cabot et al. (2021)
@@ 3.2 Template spectra for cross-correlation
The Cross-Correlation method is needed to detect the faint planetary lines in the residuals obtained after PCA. We cannot detect most of them directly except for a few individual lines (such as ). Fortunately, atoms and molecules have many spectral lines and a cross-correlation with a template boosts the signal, allowing a precise detection of the planetary spectrum. A template is needed to combine these lines. We generated synthetic spectra specific to TOI-1518 b using the parameters from Table. 2 and PetitRadtrans (Mollière et al., 2019), assuming a temperature of 2500K, with chemical abundances determined by equilibrium chemistry and calculated with FastChem (Stock et al., 2022).
We produced spectra for individual molecules at a resolution of over the 400 to 1000 nm wavelength range. These single-species spectra, to be used as cross-correlation templates, use collision induced absorption (CIA) cross-sections of H2-H2 and H2-He as continuum. We then interpolated the spectra onto the MAROON-X wavelength grid and convolved them to match the instrumental resolution.
The spectra were also convolved with the planetary rotation kernel. These spectra are shown in Fig. 3. The species selected in this study are based on those previously detected in recent MAROON-X publications (Prinoth et al., 2023; Pelletier et al., 2023).
Line list from Kurucz database were used for all atoms and ions (Kurucz, 2017). For TiO, we used TOTO line list (McKemmish et al., 2019). For VO, both the HyVO line list (Bowesman et al., 2024) and the VOmyt line list (McKemmish et al., 2016) outputs were compared in Sect.3.3 and 3.5.
Alkaline metals and ions show individual strong lines, while other metals show line forests. Most of the signals except for the triplet are stronger in the blue part of the spectrum, which corresponds to the blue detector of MAROON-X. We decided to analyze each detector individually as if it were two different transits and then to sum every - map or CCF map.
@@ 3.3 Species detected in the - maps
The cross-correlation maps are converted into velocity-velocity maps (- diagrams) by shifting them to the expected rest-frame of the planet ( = 207.32 4.51 km.s-1, = 0.17 km.s-1), assuming values of projected orbital velocity between 0 and 400 km.s-1 in steps of 1 km. s-1. Fig. 4 shows fourteen clear detections obtained via cross-correlating the signal with single-species templates. The white cross in each plot is the position of the expected signal from the planet if the planetary atmosphere is considered static and the planet has a circular orbit. The expected is calculated as follows:
with , where is the period, is the semi-major axis of TOI-1518 b and the inclination of the system. All these parameters are given in Table 2. To compute our - plots, we selected a range of orbital velocity () from 0 to 400 km.s-1 and a range of rest-frame velocity () from -150 to 150 km.s-1 with steps of 1 km.s-1 for each. We then integrate each point of the CCF maps previously obtained (see Fig. 4) following the slope determined by the orbital velocity at the rest frame position determined by . The noise level is calculated in a region far from the central peak at 75 km.s-1 where no signal of the planet or Rossiter-McLaughlin (RM) residuals is expected. Fourteen species are detected with an S/N 4. The parameters of the best Gaussian fits of the maps are presented in Table 3.
@@
| -74.1 0.5 | 21.1 0.5 | 4.9 0.1 | 36.9 0.2 | |
| -7.1 0.1 | 12.3 1.1 | -4.0 0.5 | 11.4 1.2 | |
| -16.0 0.1 | 7.5 1.1 | -3.4 0.6 | 9.1 1.4 | |
| Ba+ | -15.9 0.1 | 11.4 -0.6 | -5.2 0.3 | 13.7 0.8 |
| Ca | -11.4 0.1 | 8.1 0.9 | -3.7 0.6 | 10.1 1.4 |
| Fe | -22.2 0.2 | 17.9 1.7 | -3.8 0.5 | 11.1 1.2 |
| Ti | -4.5 0.4 | 10.1 0.7 | -4.3 0.3 | 9.7 0.8 |
| Si | -29.2 0.2 | 5.9 0.9 | -1.4 0.8 | 9.7 1.8 |
| Mg | -6.5 0.3 | 8.1 0.8 | -3.1 0.6 | 11.7 1.4 |
| V | -47.4 0.6 | 5.8 0.5 | -3.1 0.5 | 10.2 1.1 |
| Mn | -5.5 0.6 | 4.0 0.2 | -3.5 0.2 | 12.7 0.6 |
| Cr | -8.8 0.1 | 4.0 0.4 | -1.3 0.8 | 12.5 1.9 |
| Na | -14.2 0.1 | 7.9 1.0 | -1.3 1.2 | 18.3 3.3 |
| VO | -32.8 0.2 | 4.9 0.2 | -12.1 0.2 | 10.5 0.5 |
| All | -20.5 0.1 | 21.1 2.6 | -2.9 0.7 | 11.6 1.7 |
The of other species of interested detected in other UHJs are presented in Appendix (see Fig 19). These ones show positive correlation near the expected orbital position and may warrant follow-up observations. On averaged, most species have a blueshifted signal compared to the expected velocity (Vres = km.s-1). The orbital velocity is also lower than expected (km.s-1).